OpenStudy (andijo76):

i have another statistics question if anyone is interested its one of my hard questions cant figure it out

OpenStudy (tkhunny):

It's a LOT easier just to post the question and show your work.

OpenStudy (andijo76):

i know but i thought i would give it a shot

OpenStudy (andijo76):

i will post the question if anyone thinks they can help me

OpenStudy (andijo76):

i will post it anyway if someone can help wonderful if not its worth a try

OpenStudy (andijo76):

three randomly selected households are surveyed. the numbers of people in the households are 1,3,8 assume that samples of size n =2 are randomly selelcted with replacement from the population of 1,3,8. Listed below are the nine different sampls 1,1 1,3 1,8 3,1 3,3 3,8 8,1 8,3 8,8 find the variance of each of the nine samples then summerize the sampling distribution of the variances in the format of a table representing the probability distribution of the distinct variance values s^2 Probability are 3,3 0 or 1 6.3 1 or 2 12.5 6.3 or 40.5 49 12.3 or 24.5

OpenStudy (tkhunny):

The variance of each sample of 2. 1,1 Mean 1 Sum x^2 = 2 Variance 2-1 = 1 1,3 Mean 2 Sum x^2 = 10 Variance 10-2 = 8 1,8 Mean 4.5 Sum x^2 = 65 Variance 65-4.5 = 60.5 3,1 Mean 2 Sum x^2 = 10 Variance 10-2 = 8 3,3 Mean 3 Sum x^2 = 18 Variance 18-3 = 15 3,8 Mean 5.5 Sum x^2 = 73 Variance 73-5.5 = 67.5 8,1 Mean 4.5 Sum x^2 = 65 Variance 65-4.5 = 60.5 8,3 Mean 5.5 Sum x^2 = 73 Variance 73-5.5 = 67.5 8,8 Mean 8 Sum x^2 = 128 Variance 128-8 = 120 It's a little tedious, but there are shortcuts with samples of only 2.